OpenStudy (anonymous):

2) Is this right? log(2)(x+2)+log(2)(x-4)=4 x= (2+log(2) over log(2) ?????

OpenStudy (anonymous):

ok, so I way way off. I notice two division signs in there and it's throwing me off. Is this a fraction?

hero (hero):

Of course you can rewrite that in a different way

OpenStudy (anonymous):

weird, i've never seen it like that, is it simplified?

hero (hero):

It simplifies into either a whole number or decimal form

hero (hero):

log(2) (x+2 +x - 4) = 4 log(2)(2x - 2) = 4 2x - 2 = 4/log(4) x = (4/log(4) - 2)/2

hero (hero):

2.32193

OpenStudy (anonymous):

uh oh, you lost me, lol. I need an integer or simplified fraction???? I think it's the fraction I'm after.

hero (hero):

You can rewrite it as: \[x = \frac{\log 5}{\log 2}\]

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

I guess the answer was simply 6!